# Effects of Artificial Collisions, Filtering, and Nonlocal Closure Approaches on Hermite-based Vlasov-Poisson Simulations

### v1

@@ 3.4 Linear Response Function Approximation
The convergence rates of the different methods are shown in Figure 2(c). The absolute error of the response function is evaluated using the $L_{2}$ norm on a log-uniform grid of $\xi\in[10^{-2},10^{2}]$ with $10^{5}$ nodes (more nodes do not change the error significantly). The absolute error is plotted with the optimal tunable parameter of the method, such that the approximate response function matches the analytic response at the asymptotic limit of $|\xi|\ll 1$. We indicate the optimal parameters for each method in Table 3.4. For artificial collisions, the tunable parameter is the artificial collisional rate $\nu\in\mathbb{R}_{+}$ in Eq. (17). For the nonlocal closure with $N_{m}=1$, the tunable parameter is the coefficients in the linear relation $\mu_{n}\in\mathbb{C}$ in Eq. (19). For filtering, the tunable parameter is $\chi/\Delta t\in\mathbb{R}_{+}$ in Eq. (18). There is not a clear trend in optimal parameter variation as a function of the velocity resolution $N_{v}$. Interestingly, the optimal closure coefficients with $N_{m}=3$, show that $\mu_{N_{v}-2}=0$ for all considered even $N_{v}$, such that a nonlocal closure with $N_{m}=2$ and $N_{m}=1$ are equivalent. The approximate response function for both artificial collisions and filtering depends on the wavenumber $k$, i.e. $R^{aw}_{N_{v}}(\xi,k)$. We present the results with $k=1$, and additional tests (not shown here) indicate that varying $k$ does not significantly influence the convergence or accuracy of the methods. The most accurate methods are hypercollisions with $\alpha=2$ and $\alpha=3$. The hypercollisional operator with $\alpha=2$ converges about four times faster than that with $\alpha=1$ (LB collisions). Accuracy decreases when $\alpha=4$, indicating that $\alpha=2$ and $\alpha=3$ represent the optimal range for the order of artificial collisions. Additionally, the nonlocal closure method with $N_{m}=3$ converges approximately twice as fast as with $N_{m}=1$. As shown by Smith [38], the convergence rate of the nonlocal closure improves as $N_{m}$ increases. However, matching the closure coefficient becomes increasingly difficult as $N_{m}$ and $N_{v}$ increase since calculating the approximate response function requires a symbolic matrix inversion, see Eq. (15), with $N_{m}$ nonlinear solves to match to the asymptotic limit. The filtering approximation converges slowly, similar to the nonlocal closure method with $N_{m}=1$.
@@ 3.5 Linear Eigenvalue Analysis
The previous section focused on analyzing the behavior of the response function for the Vlasov equation alone. We now investigate the ability of the discrete models to replicate the dynamics of the linearized Vlasov-Poisson system, using the Landau damping problem as a test case. We set $\ell=4\pi$ such that wavenumber $k$ are integer multiples of $1/2$. The dispersion relation $\omega(k)$ for linear Landau damping is given by $k^{2}=R(\xi)$. We adopt a more straightforward approach [1, 20, 2] by numerically calculating the damping rate through the real component of the least damped eigenvalue of the discretized linear Vlasov-Poisson equations:
@@
Figure 4 shows the relative error of the Landau damping rate for various spatial wavenumbers as a function of the system’s tunable parameter with $N_{v}=20$. The nonlocal closure underestimates the damping rate for all spatial wavenumbers, irrespective of the value of the tunable parameter, so that higher wavenumber modes are most severely underdamped. Furthermore, no value of the closure coefficient $\mu_{N_{v}-1}$ can spectrally alter the discrete system to accurately recover the correct damping rate. The same underdamping effect also applies to the filtering technique. The LB operator, i.e. artificial collisions with $\alpha=1$, can replicate the correct damping for a single spatial mode, but not for a range of wavenumbers. As a result, if the system is initially excited with a range of modes, accurately capturing the damping rate of the highest wavenumber results in the LB operator overdamping all other modes, which is consistent with the results in Pezzi et al. [20]. In contrast, setting $\alpha=2$ (as in Camporeale et al. [23]) does not overdamp modes. As we increase $\alpha$ to $3$ and $4$, the higher wavenumber modes are underdamped, such that $\alpha=2$ provides the ideal balance.
@@ 4 Numerical Results
We use the AW Hermite discretization in velocity and Fourier discretization in space, as described in [23, 15], to simulate Landau damping with artificial collisions, filtering, and nonlocal closure approaches discussed in section 3. The linear and nonlinear Landau damping tests are initialized by perturbing the initial electron distribution function as follows
@@ 4.1 Linear Landau Damping
We investigate the impact of artificial collisions, filtering, and nonlocal closure approaches on the recurrence problem in linear Landau damping with $N_{v}=20$. Table 4 lists the optimal tunable parameters for each method, which are derived through the eigenvalue analysis in section 3.5. Initially, two modes with $k=0.5$ and $k=1.5$ are perturbed with amplitude $\epsilon=10^{-2}$ in Eq. (20). The velocity space Hermite moment cascade is shown in Figure 6. We plot the squared normalized magnitude of the Hermite coefficients with $k=0.5$, more specifically $|\hat{C}_{n}(k=0.5,t)|^{2}/\max_{n}(|\hat{C}_{n}(k=0.5,t)|^{2})$ [43]. The results show transport from the zeroth to the highest order Hermite mode, which for the collisionless, nonlocal closure, and filtering, gets reflected from the finite Hermite resolution and backpropagates. This backpropagation then artificially amplifies the zeroth order Hermite mode, resulting in a growth of the electric field. Artificial collisions suppress higher-order Hermite modes and prevent artificial backpropagation of Hermite flux.
@@
Figure 6: The squared normalized magnitude of the Hermite coefficients with $k=0.5$ for $N_{v}=20$. The results indicate that artificial collisions in subfigures (c)–(e) effectively mitigate recurrence caused by an artificial backpropagating Hermite cascade due to truncation, whereas nonlocal closure in subfigure (b) and filtering in subfigure (f) do not adequately suppress this backpropagation.
@@
Figure 7: The linear Landau damping electric field Fourier coefficient magnitude of $k\in\{0.5,1.5\}$ with $N_{v}=20$. The solid lines show the analytic damping rate obtained from linear theory, i.e. solving $k^{2}=R(\xi)$. Recurrence time is given by $T_{\mathrm{rec}}(k)\approx\sqrt{N_{v}}/k$, such that recurrence occurs at $T_{\mathrm{rec}}(k=1.5)\approx 4$ and $T_{\mathrm{rec}}(k=0.5)\approx 13$, as shown in subfigure (a). The numerical results show that only artificial collisions with $\alpha=2$ in subfigure (c) can mitigate recurrence without overdamping the smallest wavenumber perturbations or underdamping the largest wavenumber perturbations.
4.2 Nonlinear Landau Damping
The electron distribution function in phase space is shown in Figure 10, which becomes highly oscillatory due to filamentation in velocity space. The oscillations then roll up to form a vortex centered at the phase speed $\omega_{r}/k=\pm 2.8$. The velocity resolution of $N_{v}=300$ does not resolve the filamentation microstructures at $t=40$, yet all methods besides LB collisions capture the correct particle trapping vortex around the phase speed. The phase space dynamics are distorted by numerical filamentation artifacts in the collisionless case with closure by truncation. The filamentation artifacts shown at $t=40$ for the nonlocal closure and filtering approaches are a result of limited velocity resolution. The hypercollisions simulation with $\alpha\geq 2$ smooths out any strong deformations of the particle distribution function while maintaining the correct plasma dynamics characteristics. Additionally, Figure 11 shows a cross-section of the electron distribution function at $t=40$ and $x=3\pi$, highlighting that artificial collisions are more effective in preserving the positivity of the electron distribution function compared to nonlocal closure and filtering techniques. Pagliantini et al. [12] showed that positivity can be improved by dynamically adjusting the Hermite tunable parameters over time or increasing the artificial collisional frequency $\nu$ in Eq. (17) with $\alpha=2$. This implies that the loss of positivity arises from insufficient spectral resolution. It is worth noting that even the reference solution exhibits negativity near $v\approx 2.5$, as the finite difference scheme does not enforce positivity.
(a) Collisionless with
closure by truncation
(b) Nonlocal closure with $N_{m}=1$
Smith [38]
(c) Artificial collisions $\alpha=1$
Lenard and Bernstein [29]
(d) Artificial collisions $\alpha=2$
Camporeale et al. [23]
(e) Artificial collisions $\alpha=3$
(f) Hou and Li [31] filter
Figure 8: Same as Figure 6 for nonlinear Landau damping with $N_{v}=300$. The nonlinear dynamics lead to the backward propagation of Hermite flux from higher to lower moments. Moreover, the finite dimensionality in velocity space introduces an artificial backward propagation of Hermite flux. The results from the nonlocal closure in subfigure (b) show that the method mitigates some of this backward propagation compared to subfigure (a). Artificial collisions in subfigures (c)–(e) and filtering in subfigure (g) significantly damp the highest modes.
@@ 5 Conclusions
The analysis highlights hypercollisions as a practical approach for enhancing accuracy in kinetic plasma simulations, as they effectively suppress artificial recurrence without compromising the accurate representation of collisionless dynamics. Future research could explore extending these methods to electromagnetic Vlasov-Maxwell simulations, including deriving nonlocal closures in Hermite space that capture electromagnetic wave-particle interaction such as cyclotron resonance similar to the four-moment fluid model in Jikei and Amano [57] but generalized to higher-moment models in Hermite space. Additionally, it would be interesting to examine a hybrid formulation of the different methods, e.g. nonlocal closure method coupled with filtering or artificial collisions.
@@ Appendix A Klimas Filter
$\bar{f}(x,v,t)\coloneqq\Theta(v)\ast f(x,v,t)=\int_{\mathbb{R}}\Theta(v-v^{\prime})f(x,v^{\prime},t)\mathrm{d}v^{\prime},\qquad\mathrm{where}\qquad\Theta(v)\coloneqq\frac{1}{\sqrt{2\pi}v_{0}}\exp\left(-\frac{1}{2}\left(\frac{v}{v_{0}}\right)^{2}\right).$
@@
The filter affects only the right-hand side of the Vlasov equation, introducing a second derivative that combines velocity and spatial derivatives. Additionally, the charge density is invariant under filtering $\int_{\mathbb{R}}f(x,v,t)\mathrm{d}v=\int_{\mathbb{R}}\bar{f}(x,v,t)\mathrm{d}v$, such that the evolution of the electric field remains unchanged [32]. By applying the AW Hermite discretization in velocity, Fourier discretization in space, and using the recursive property in Eq. (11), we obtain
$\mathcal{C}_{\mathrm{Klimas}}(\bar{f})\coloneqq-v_{0}^{2}\frac{\partial^{2}}{\partial x\partial v}\bar{f}\qquad\Rightarrow\qquad\mathcal{C}_{\mathrm{Klimas}}(\hat{C}_{n})=ikv_{0}^{2}\sqrt{n}\hat{C}_{n-1}.$
We repeat the eigenvalue analysis in section 3.5 and find that the Klimas filter cannot recover damping, i.e. $\gamma<0$. Figure 12 shows that the Klimas filter predicts that higher wavenumber modes have an instability or no energy exchange between the wave and the particle, i.e. $\gamma\geq 0$, regardless of $v_{0}\in\mathbb{R}$. Note that it is important to maintain $v_{0}\lessapprox 1$ since the convolution of two Gaussian distributions is $\mathcal{G}\left(\mu_{1},\sigma_{1}\right)\ast\mathcal{G}\left(\mu_{2},\sigma_{2}\right)=\mathcal{G}\left(\mu_{1}+\mu_{2},\sqrt{\sigma_{1}+\sigma_{2}}\right)$, such that setting the parameter $v_{0}$ much greater than the electron thermal velocity results in oversmoothing of the solution and thus incorrect instability. Thus, this method is not viable for recovering the correct Landau root and accurately capturing the long-term behavior resulting from phase mixing generated by the continuous spectrum in AW Hermite-based solvers. Klimas and Viñas [32] show that the Klimas filter can suppress recurrence in doubly Fourier-transformed solvers with imposed outflow boundary conditions; however, our analysis shows that it fails for Hermite discretization in velocity space with closure by truncation [58].
(a) $N_{v}=20$
(b) $N_{v}=100$
Figure 12: Landau damping rate relative error of the Klimas [30] filter approach denoted by $\gamma^{*}$ with (a) $N_{v}=20$ and (b) $N_{v}=100$ compared with analytic damping rate denoted by $\gamma$. Both velocity resolutions are not able to recover damping and instead predict no exchange of energy between the particle and the electrostatic wave or an instability ($\gamma\geq 0$) for all values of the free parameter $v_{0}\in\mathbb{R}$. The Klimas filter is not viable for recovering the correct Landau root in Hermite-based simulations.

### v2

@@ 3.4 Linear Response Function Approximation
The convergence rates of the different methods are shown in Figure 2(c). The absolute error of the response function is evaluated using the $L_{2}$ norm on a log-uniform grid of $\xi\in[10^{-2},10^{2}]$ with $10^{5}$ nodes (more nodes do not change the error significantly). The absolute error is plotted with the optimal tunable parameter of the method, such that the approximate response function matches the analytic response at the asymptotic limit of $|\xi|\ll 1$. We indicate the optimal parameters for each method in Table 3.4. For artificial collisions, the tunable parameter is the artificial collisional rate $\nu\in\mathbb{R}_{+}$ in Eq. (19). For the nonlocal closure with $N_{m}=1$, the tunable parameter is the coefficient in the linear relation $\mu_{n}\in\mathbb{C}$ in Eq. (21). For filtering, the tunable parameter is $\chi/\Delta t\in\mathbb{R}_{+}$ in Eq. (20). There is not a clear trend in optimal parameter variation as a function of the velocity resolution $N_{v}$. Interestingly, the optimal closure coefficients with $N_{m}=3$, show that $\mu_{N_{v}-2}=0$ for all considered even $N_{v}$, such that a nonlocal closure with $N_{m}=2$ and $N_{m}=1$ are equivalent. The approximate response function for both artificial collisions and filtering depends on the wavenumber $k$, i.e. $R^{aw}_{N_{v}}(\xi,k)$. We present the results with $k=1$, and additional tests (not shown here) indicate that varying $k$ does not significantly influence the convergence or accuracy of the methods. The most accurate methods are hypercollisions with $\alpha=2$ and $\alpha=3$. The hypercollisional operator with $\alpha=2$ converges about four times faster than that with $\alpha=1$ (LB collisions). Accuracy decreases when $\alpha=4$, indicating that $\alpha=2$ and $\alpha=3$ represent the optimal range for the order of artificial collisions. Additionally, the nonlocal closure method with $N_{m}=3$ converges approximately twice as fast as with $N_{m}=1$. As shown by Smith [38], the convergence rate of the nonlocal closure improves as $N_{m}$ increases. However, matching the closure coefficient becomes increasingly difficult as $N_{m}$ and $N_{v}$ increase since calculating the approximate response function requires a symbolic matrix inversion, see Eq. (17), with $N_{m}$ nonlinear solves to match to the asymptotic limit. The filtering approximation converges slowly, similar to the nonlocal closure method with $N_{m}=1$.
@@ 3.5 Linear Eigenvalue Analysis
The previous section focused on analyzing the behavior of the response function for the Vlasov equation alone. We now investigate the ability of the discrete models to replicate the dynamics of the linearized Vlasov-Poisson system, using the Landau damping problem as a test case. We set $\ell=4\pi$ such that wavenumber $k$ are integer multiples of $1/2$. The dispersion relation $\omega(k)$ for linear Landau damping is given by $k^{2}=-R(\xi)$. We adopt a more straightforward approach [1, 20, 2] by numerically calculating the damping rate through the real component of the least damped eigenvalue of the discretized linear Vlasov-Poisson equations:
@@
Figure 4 shows the relative error of the Landau damping rate for various spatial wavenumbers as a function of the system’s tunable parameter with $N_{v}=20$. In this and the following sections, we restrict the nonlocal closure to $N_{m}=1$ to avoid higher-dimensional parametric optimization and to constrain each method to a single free parameter. The nonlocal closure underestimates the damping rate for all spatial wavenumbers, irrespective of the value of the tunable parameter, so that higher wavenumber modes are most severely underdamped. Furthermore, no value of the closure coefficient $\mu_{N_{v}-1}$ can spectrally alter the discrete system to accurately recover the correct damping rate. The same underdamping effect also applies to the filtering technique. The LB operator, i.e. artificial collisions with $\alpha=1$, can replicate the correct damping for a single spatial mode, but not for a range of wavenumbers. As a result, if the system is initially excited with a range of modes, accurately capturing the damping rate of the highest wavenumber results in the LB operator overdamping all other modes, which is consistent with the results in Pezzi et al. [20]. In contrast, setting $\alpha=2$ (as in Camporeale et al. [23]) does not overdamp modes. As we increase $\alpha$ to $3$ and $4$, the higher wavenumber modes are underdamped, such that $\alpha=2$ provides the ideal balance.
@@ 4 Numerical Results
We use the AW Hermite discretization in velocity and Fourier discretization in space, as described in Camporeale et al. [23], to simulate Landau damping with artificial collisions, filtering, and nonlocal closure approaches discussed in section 3. The linear and nonlinear Landau damping tests are initialized by perturbing the initial electron distribution function as follows
@@ 4.1 Linear Landau Damping
We investigate the impact of artificial collisions, filtering, and nonlocal closure approaches on the recurrence problem in linear Landau damping with $N_{v}=20$. Table 4 lists the optimal tunable parameters for each method, which are derived through the eigenvalue analysis in section 3.5. Initially, two modes with $k=0.5$ and $k=1.5$ are perturbed with amplitude $\epsilon=10^{-2}$ in Eq. (22). The velocity space Hermite moment cascade is shown in Figure 6. We plot the squared normalized magnitude of the Hermite coefficients with $k=0.5$, more specifically $|\hat{C}_{n}(k=0.5,t)|^{2}/\max_{n}(|\hat{C}_{n}(k=0.5,t)|^{2})$. The results show transport from the zeroth to the highest order Hermite mode, which for the collisionless, nonlocal closure, and filtering, gets reflected from the finite Hermite resolution and backpropagates. For the linear Landau damping problem, Hermite flux only propagates from the lower to higher order Hermite modes as shown in Figure 5(a) for the collisionless simulation with closure by truncation and $N_{v}=1,024$, and explained in greater detail in Parker and Dellar [43]. This backpropagation then artificially amplifies the zeroth order Hermite mode, resulting in a growth of the electric field. Artificial collisions suppress higher-order Hermite modes and prevent artificial backpropagation of Hermite flux.
@@
(a) Collisionless with
closure by truncation
$N_{v}=1,024$
(b) Collisionless with
closure by truncation
$N_{v}=20$
(c) Nonlocal closure with $N_{m}=1$
Smith [38]
(d) Artificial collisions $\alpha=1$
Lenard and Bernstein [29]
(e) Artificial collisions $\alpha=2$
Camporeale et al. [23]
(f) Artificial collisions $\alpha=3$
(g) Hou and Li [31] filter
Figure 6: The squared normalized magnitude of the Hermite coefficients with $k=0.5$ for (a) $N_{v}=1,024$ and (b)-(g) $N_{v}=20$. The results indicate that artificial collisions in subfigures (d)–(f) effectively mitigate recurrence caused by an artificial backpropagating Hermite cascade due to truncation, whereas nonlocal closure in subfigure (c) and filtering in subfigure (g) do not adequately suppress this backpropagation.
@@
Figure 7: The linear Landau damping electric field Fourier coefficient magnitude of $k\in\{0.5,1.5\}$ with $N_{v}=20$. The solid lines show the analytic damping rate obtained from linear theory, i.e. solving $k^{2}=-R(\xi)$. Recurrence time is given by $T_{\mathrm{rec}}(k)\approx\sqrt{N_{v}}/k$, such that recurrence occurs at $T_{\mathrm{rec}}(k=1.5)\approx 4$ and $T_{\mathrm{rec}}(k=0.5)\approx 13$, as shown in subfigure (a). The numerical results show that only artificial collisions with $\alpha=2$ in subfigure (c) can mitigate recurrence without overdamping the smallest wavenumber perturbations or underdamping the largest wavenumber perturbations.
4.2 Nonlinear Landau Damping
The electron distribution function in phase space is shown in Figure 10, which becomes highly oscillatory due to filamentation in velocity space. The oscillations then roll up to form a vortex centered at the phase speed $\omega_{r}/k=\pm 2.8$. The velocity resolution of $N_{v}=300$ does not resolve the filamentation microstructures at $t=40$ by comparison to the reference solution in Figure 9(g), yet all methods besides LB collisions capture the correct particle trapping vortex around the phase speed. At $t=40$, the shortest velocity wavelength of the filaments in the reference solution is $\lambda_{v}\approx 0.01$. According to Eq. (10), resolving such velocity scales with the Hermite discretization requires $N_{v}\approx(2\pi/\lambda_{v})^{2}=394,784$. The phase space dynamics are distorted by numerical filamentation artifacts in the collisionless case with closure by truncation. The filamentation artifacts shown at $t=40$ for the nonlocal closure and filtering approaches are a result of limited velocity resolution. The hypercollisions simulation with $\alpha\geq 2$ smooths out any strong deformations of the particle distribution function while maintaining the correct plasma dynamics characteristics. Additionally, Figure 11 shows a cross-section of the electron distribution function at $t=40$ and $x=3\pi$, highlighting that artificial collisions are more effective in preserving the positivity of the electron distribution function compared to nonlocal closure and filtering techniques. In the spectral solver, the particle distribution function is not guaranteed to remain positive due to the orthogonal nature of the velocity basis functions, potentially leading to unphysical solutions. To enforce positivity, a square-root formulation can be used; see the work by Issan et al. [10], which employs this transformation for the symmetrically weighted Hermite discretization. Pagliantini et al. [12] showed that positivity can be improved by dynamically adjusting the Hermite tunable parameters over time or increasing the artificial collisional frequency $\nu$ in Eq. (19) with $\alpha=2$. This implies that the loss of positivity arises from insufficient spectral resolution. It is worth noting that even the reference solution exhibits negativity near $v\approx 2.5$, as the finite difference scheme does not enforce positivity.
@@ 4.2 Nonlinear Landau Damping
Figure 8: Same as Figure 6 for nonlinear Landau damping with $N_{v}=300$. The nonlinear dynamics lead to the backward propagation of Hermite flux from higher to lower moments. Moreover, the finite dimensionality in velocity space introduces an artificial backward propagation of Hermite flux. The results from the nonlocal closure in subfigure (b) show that the method mitigates some of this backward propagation compared to subfigure (a). Artificial collisions in subfigures (c)–(e) and filtering in subfigure (f) significantly damp the highest modes.
@@ 5 Conclusions
The analysis highlights hypercollisions as a practical approach for enhancing accuracy in kinetic plasma simulations, as they effectively suppress artificial recurrence without compromising the accurate representation of collisionless dynamics. Future research could explore extending these methods to electromagnetic Vlasov-Maxwell simulations, including deriving nonlocal closures in Hermite space that capture electromagnetic wave-particle interaction such as cyclotron resonance similar to the four-moment fluid model in Jikei and Amano [57] but generalized to higher-moment models in Hermite space. Additionally, it would be interesting to examine a hybrid formulation of the different methods, e.g. nonlocal closure method coupled with filtering or artificial collisions. Lastly, future work could explore the response function approximation for non-Maxwellian equilibrium distributions [58], such as Kappa or Cauchy distributions with suprathermal tails [59]. For these non-Maxwellian distributions, the finite AW Hermite expansion would approximate both the equilibrium and perturbation components, whereas the zeroth term in the AW Hermite expansion represents exactly a Maxwellian equilibrium. Additionally, other spectral discretizations, such as the Legendre basis, may be well-suited for non-Maxwellian distributions, which similarly to the AW Hermite basis also exhibit fluid-kinetic coupling properties [60].
@@ Appendix A Klimas Filter
$\bar{f}(x,v,t)\coloneqq\int_{\mathbb{R}}\Theta(v-v^{\prime})f(x,v^{\prime},t)\mathrm{d}v^{\prime},\qquad\mathrm{where}\qquad\Theta(v)\coloneqq\frac{1}{\sqrt{2\pi}v_{0}}\exp\left(-\frac{1}{2}\left(\frac{v}{v_{0}}\right)^{2}\right).$
@@
The filter modifies only the right-hand side of the Vlasov equation, introducing a second derivative that combines velocity and spatial derivatives. Additionally, the charge density and current density are invariant under filtering, i.e. $\int_{\mathbb{R}}f(x,v,t)\mathrm{d}v=\int_{\mathbb{R}}\bar{f}(x,v,t)\mathrm{d}v$ and $\int_{\mathbb{R}}vf(x,v,t)\mathrm{d}v=\int_{\mathbb{R}}v\bar{f}(x,v,t)\mathrm{d}v$, such that the evolution of the electric field remains unchanged for any value of $v_{0}\in\mathbb{R}$ [30]. By applying the AW Hermite discretization in velocity, Fourier discretization in space, and using the recursive property in Eq. (12), we obtain
$-v_{0}^{2}\frac{\partial^{2}}{\partial x\partial v}\bar{f}\qquad\Rightarrow\qquad ikv_{0}^{2}\sqrt{n}\hat{C}_{n-1}.$
Because the Klimas filter does not correspond to a true dissipation, the Landau root is not represented by a discrete mode as in the cases considered in the main text. We have therefore tested it with dynamical simulations with different values of $v_{0}$ and found that recurrence effects are not eliminated when the Klimas filter is applied to the AW Hermite discretization with closure by truncation. Indeed,  Klimas and Viñas [32] show that the Klimas filter can suppress recurrence in doubly Fourier-transformed solvers with imposed proper boundary conditions. Note that it is important to maintain $v_{0}\lessapprox 1$ since the convolution of two Gaussian distributions is $\mathcal{G}\left(\mu_{1},\sigma_{1}\right)\ast\mathcal{G}\left(\mu_{2},\sigma_{2}\right)=\mathcal{G}\left(\mu_{1}+\mu_{2},\sqrt{\sigma_{1}+\sigma_{2}}\right)$, such that setting the parameter $v_{0}$ much greater than the electron thermal velocity results in oversmoothing of the solution and the same eigenvalue analysis performed in section 3.5 predicts discrete unstable modes.